OpenStudy (anonymous):

What is the perimeter of a triangle with the given side lengths? s1 = x – 3 ft s2 = x + 2 ft s3 = x + 4 ft P = x + ft

OpenStudy (bibby):

the perimeter is the summation of all the sides. \(s_1+s_2+s_3=?\)

OpenStudy (anonymous):

6s

OpenStudy (camerondoherty):

The perimeter is the sum of all the sides, correct? So just add the sides together \[p =(x-3)+(x+2)+(x+4)\] The solution is the perimeter, which is the sum of all sides

OpenStudy (bibby):

\(s_1=x-3\) for example, \(s_1+3=x-3+3=x\)

OpenStudy (anonymous):

so what is the answer

OpenStudy (camerondoherty):

Solve the equation I gave you first, and you'll get it

OpenStudy (anonymous):

9x

OpenStudy (camerondoherty):

Are you sure that comes out to 9x?

OpenStudy (camerondoherty):

Or did you just take a guess?

OpenStudy (jdoe0001):

*cough*

OpenStudy (bibby):

hmm

OpenStudy (jdoe0001):

perimeter = \(s_1+s_2+s_3\) so add them up as camerondoherty suggested

OpenStudy (jdoe0001):

think about it why did you teacher give you the answer sheet along with the exercises? if he/she just wanted you to transcribe them over, he/she would have handed out the answer sheet as well but that defeats the idea behind doing the exercise at all

OpenStudy (anonymous):

@abswag13